# Non-Adiabatic Quantum Optimization for Crossing Quantum Phase Transitions

### v2

@@ III Optimal control based on the adiabatic condition
In the next sections, we inspect the performance of the two approaches outlined above in the TFIM: the LAD is treated in Sec. III.1 and the QAB in Sec. III.2. Taking inspiration from the KZ mechanism, we look at the scaling of the defect number with the final time $T$ when the schedule $g(t)$ is not a linear function but instead determined by Eqs. (15) or (16). We also compute the timescale of the onset of adiabaticity, i.e.,  the time $T_{\mathrm{ons}}$, after which the defect number becomes of $\mathcal{O}(1)$. This timescale is not to be confused with $T_{\mathrm{ad}}$, which always refers to the global adiabatic condition, Eq. (14). In Sec. III.3, we prove a simple connection between the two schemes, LAD and QAB, that shows how the two are interchangeable under some approximation.
III.1 Local adiabatic driving
@@
III.2 Quantum adiabatic brachistochrone
In the previous section, it was argued that the LAD procedure fails to produce sensible annealing schedules $g(t)$. Slowing down at the avoided crossing of the zero mode can only be achieved at the expense of speeding up near the end of the evolution, exciting an extensive number of momentum modes. The quantum adiabatic brachistochrone (QAB) approach tries to solve this problem by weighting the adiabatic condition along all the paths through the action $S_{\mathrm{QAB}}$, Eq. (16), rather than applying it locally. When specified to the TFIM, the action reads
@@ III.2 Quantum adiabatic brachistochrone
where again $k_{0}=\pi/L$ is the lowest momentum for a system of size $L$. Notice that the norm $\|\partial_{t}\hat{H}\|^{2}=L\,\dot{g}^{2}$, as one can check in the fermionic basis (see Sec. II.2). The Euler-Lagrange equation for the action reads
@@
In order to investigate the efficiency of the QAB schedules, the scaling of the number of defects was computed numerically, again taking inspiration from the KZ studies. From Fig. 3b one can see that the defect density remains extensive up to a time $T_{\text{plateau}}\sim L$; then, a rapid decay takes place, with $n\sim T^{-2}$. It follows that the onset of adiabaticity, defined by $n(T_{\mathrm{ons}})\approx 1/L$, takes a time $T_{\mathrm{ons}}\sim L^{3/2}$. The QAB approach does not provide efficient schedules $g_{\mathrm{QAB}}(t)$ due to the shortcoming of the zero mode, as in the LAD approach. Even though the zero mode can be driven adiabatically in a timescale $T\sim L$, this comes at the price of creating an extensive number of excitations at the avoided level crossing of higher momentum modes. Similarly to the LAD approach, the universal KZ regime does not emerge during the whole process.
@@ IV Optimal control beyond the adiabatic condition
In Sec. III, we considered optimal control strategies that take as starting point the adiabatic condition, i.e., that require the drive to be slow with respect to the inverse gap and transition matrix elements, Eq. (14). In this Section, we introduce optimal control methods that go beyond the adiabatic condition, taking into account also the higher excited states. Specifically, in Sec. IV.1, we extend the QAB to tame also transitions in excited states, while in Sec. IV.2, we introduce a new action to be minimized, based on the avoided level crossing structure in the LZ formula. Finally, we compare our results with other well-known optimization schemes: the ONLP through a quantum critical point in Sec. IV.3 and the fully numerical CRAB optimization in Sec. IV.4.
@@ IV.1 Extended quantum adiabatic brachistrochrone
where |$\ket{\epsilon_{\pm,k}(t)}$ denotes the ground (--) and excited ($+$) single-particle states in the $k$-th momentum mode, and $2\epsilon_{k}(t)$ equals the gap in the same mode. We call this action “extended quantum adiabatic brachistochrone” (eQAB). Notice that in the numerator of Eq. (34) there is the matrix element between the one-particle ground and excited states, rather than the norm $\|\partial_{t}\hat{H}\|$, as described in Sec. III.3.
@@
The solution $g_{\mathrm{eQAB}}(t)$ of Eq. (40) is plotted in Fig. 7. Interestingly, the typical shape of standard and extended QAB schedules are of the opposite nature. The former is extremely flat around $g=1$, with a slope scaling as $\sim(TL)^{-1}$, and presents a sharp jump around $t=\pm T/2$. The latter is more similar to the linear schedule, without dramatic changes in the local velocity $\dot{g}(t)$.
@@
The reason why the eQAB approach does not provide a significant improvement over the QAB can be understood as follows. Equation (34) is an unweighted average over all momentum modes of an action of the QAB form. There are two consequences to this fact. First, the zero-momentum mode still dominates, as exemplified by the Euler-Lagrange equation becoming singular in the thermodynamic limit. Second, the main contribution to the excitations arises from the vanishingly small regions around the avoided crossings. However, the action in Eq. (34) does not weight properly these short time intervals where the gaps are closing, masking their true impact. These two facts lead to a considerably weaker effective condition to suppress excitations.
@@ IV.2 Non-Adiabatic Quantum Optimization (NAQO) of Landau-Zener transitions
Above, it was argued that tailoring control schedules by “extending” the adiabatic condition, Eq. (14), leads to suboptimal results, as higher-excited modes are simply not considered. Remarkably, no relevant improvement in the timescale of the onset of adiabaticity was found even by the simultaneous optimization of all the momentum modes with the eQAB, see Sec. IV.1. These results indicate that even in simple cases, many-body effects lead to non-trivial behavior, deteriorating the efficiency of the conventional optimal control strategies.
In this section, we show that it is possible to design better schedules by abandoning the adiabatic condition in favor of the LZ formula. As anticipated at the end of the last section, the schedule $g(t)$ needs to go slow only when there is a high probability of exciting a momentum mode, i.e., near an avoided crossing. This is exactly where the LZ formula applies, both in the KZ scaling regime and beyond the onset of adiabaticity [12, 78, 82, 83, 13]. Designing efficient control schedules requires the simultaneous suppression of LZ transitions in both low and high-momentum states. To capture this in an analytical way, the following action is introduced:
@@
where $g_{c}(k):=g(t_{c}(k))$ is the position of the avoided crossing as a function of $k$; see Eq. (10). We denote Eq. (41) as the NAQO action. Notice that $S_{\mathrm{NAQO}}$ is defined only for values of $g<1$, i.e., where there actually are avoided crossings. How to extend the action for $g>1$ is explained below.
@@ IV.4 Fully numerical optimization
Figure 6: Comparison of the analytical NAQO and ONLP with the average and best CRAB numerical optimization. The system size is fixed with $L=1000$ since the CRAB algorithm becomes numerically demanding for larger systems. While, on average, the CRAB search does not find good schedules, the best instance over $1000$ trials falls near the ONLP results. The NAQO results shows an anomalously fast decay here due to finite-size effects, while they approach $n\sim T^{-1}$ at larger sizes, see Fig. 4.
@@ V.2 Optimal control in the universal KZ scaling regime
In this regime, the behavior of the low momentum modes slightly changes, and a more pronounced role goes to their suppression, as entailed by $\dot{g}^{3}/(2-g)^{(2\beta+\alpha-3)/(\alpha-1)}$. The second term, however, only acquires an $\alpha$ and $\beta$ dependent prefactor, accounting in the same way for the higher momentum excitations. Here, the $\alpha$ independence is only observed for the linear and ONLP schedules, while increasing $\alpha$ shifts the LZ-optimized schedules by a constant factor, as observed in Fig. 12. This feature is attributed to the additional $\alpha$-dependent exponent in the corresponding differential equation. Despite these additional features, the resulting NAQO schedule provides the fastest decay of the defects for all $\alpha<3$.
@@
where again $W$ is the Lambert function, cf. Eq. (46). At small $t$ values, the denominator can be neglected, while for intermediate values, the shape stays close to the linear one as well.
Note that this limiting differential equation is of completely opposite nature to that in the eQAB, Eq. (40), which only involves the $\dot{g}^{2}/(2-g)$ term.
@@ V.3 Optimal control in the dynamical power-law-scaling regime
$4g(0)=\sin(\alpha\pi/2)\frac{\Gamma(1-\alpha)}{\zeta(\alpha)}k^{\alpha-1},$ (77)
where notably the low-momentum term on the r.h.s., $\dot{g}^{3}/(1-g)^{\frac{\alpha+1}{\alpha-1}}$, only depends now on $\alpha$. In this regime, both ONLP and NAQO provide a faster decay of $n(T)$ as $\alpha$ increases, while the linear schedule results remains unchanged according to Eq. (75), as predicted in Ref. [82].
@@ VI Demonstration on a non-integrable disordered many-body system
The model in Eq. (78) does not admit a separation into independent momentum modes, but instead, the Hilbert space is composed of only two $2^{L-1}$-dimensional sectors, according to the parity $\prod_{i}\hat{\sigma}_{i}^{x}$. We concentrate on the even parity sector, which is the one dynamically accessible from the paramagnetic ground state. While excitations continue to be generated mainly via LZ transitions at multiple avoided crossings, these transitions are no longer independent of one another. However, for the numerically achievable system sizes $L\lesssim 20$, the ground-state energy gap approximately decreases as $\sim L^{-0.65}$ (see inset of Fig. 15), and this is the most relevant LZ transition to tame. Meanwhile, avoided crossings of higher excited states appear with gradually increasing gaps and at smaller values of $g(t)$ motivating the framework of the NAQO. For this reason, a feasible ansatz for the optimal control schedule is provided by the LRKM differential equation in the long-range hopping and short-range pairing sector, $\beta<2,\,\alpha>3$ (see Sec. V)
@@
In the first term on the r.h.s., the $L$ dependence of the gap at the first avoided level crossing is captured by $\beta$. The second term accounts for the suppression of LZ transitions to higher excited states. For the numerically reachable system sizes, $L\lesssim 20$, the effective system size dependence of the ground state gap sets the exponent based on the dispersion relation of the LRMs, $\beta=1.65$, while the second term on the r.h.s. provides a good approximation for the higher excited states as it depends on $\beta$ only through a multiplicative factor. Additionally, the limit $\alpha\rightarrow\infty$ was taken, as it provides a more compact ansatz. The initial condition of the corresponding optimal differential equation is identified with the position of the gap to the ground state, also found numerically. Thus, the corresponding optimal control equation reads
@@
The results are shown in Fig. 15 for $L=14$. The present method provides a tangible speed-up compared to earlier path optimization methods [108, 109, 110], which used purely numerical optimization and typically require also the use of intermediate control Hamiltonians. In these studies, both the computational resources and the achievable efficiency are much larger than the semi-analytical NAQO method.
@@ VII Conclusions and perspectives
A challenging and promising prospect is to apply NAQO and the LZ formula to non-integrable models and in more general quantum annealing contexts. While we have demonstrated for the specific example of Sec. VI that NAQO does provide some speed-up at small system sizes, it is an open question whether NAQO is able to speed up the annealing also when, at larger sizes, the scaling of the gap becomes exponential in the system size [12]. The same exponential scaling is at the base of the difficulty of applying quantum annealing to NP-hard problems [105, 106]. In this context, optimal control is still largely uncharted, and the few existing studies point to the presence of interesting phases and controllability transitions [111, 112, 113]. Another interesting direction would be to use the analytically-optimized NAQO schedules as a starting point for further numerical optimization in the spirit of Refs. [58, 114].

### v3

@@ III Optimal control based on the adiabatic condition
In the next sections, we inspect the performance of the two approaches outlined above in the TFIM: the LAD is treated in Sec. III.1 and the QAB in Sec. III.2. Taking inspiration from the KZ mechanism, we look at the scaling of the defect number with the final time $T$ when the schedule $g(t)$ is not a linear function but instead determined by Eqs. (15) or (16). We also compute the timescale of the onset of adiabaticity, i.e.,  the time $T_{\mathrm{ons}}$, after which the defect number becomes of $\mathcal{O}(1)$. This timescale is not to be confused with $T_{\mathrm{ad}}$, which always refers to the global adiabatic condition, Eq. (14).
In short, both approaches fail to suppress excitations efficiently as they are governed by the adiabatic condition of the zero mode. From a broader perspective, this is the consequence of the underlying many-body structure, allowing for a large variety of pathways to excite the system. In Sec. III.3, we prove a simple connection between the two schemes, LAD and QAB, that shows how the two are interchangeable under some approximation.
III.1 Local adiabatic driving (LAD)
@@
III.2 Quantum adiabatic brachistochrone (QAB)
In the previous section, it was argued that the LAD procedure fails to produce sensible annealing schedules $g(t)$. Slowing down at the avoided crossing of the zero mode can only be achieved at the expense of speeding up near the end of the evolution, exciting an extensive number of momentum modes. The QAB approach tries to solve this problem by weighting the adiabatic condition along all the paths with the norm of the derivative of the Hamiltonian, i.e.,  using the action $S_{\mathrm{QAB}}$, Eq. (16), rather than applying it locally. When specified to the TFIM, the action reads
@@ III.2 Quantum adiabatic brachistochrone (QAB)
where again $k_{0}=\pi/L$ is the lowest momentum for a system of size $L$. Notice that the norm $\|\partial_{t}\hat{H}\|^{2}=L\,\dot{g}^{2}$, as one can check in the fermionic basis (see Sec. II.2). Despite equal weighting via the norm, the final expression is still governed by the zero mode energy gap. Therefore, the behavior is expected to be similar to that in the LAD approach discussed in Sec. III.1.
The Euler-Lagrange equation for the action reads
@@
In order to investigate the efficiency of the QAB schedules, the scaling of the number of defects was computed numerically, again taking inspiration from the KZ studies. In agreement with the structure of the action in Eq. (34), Fig. 3b demonstrates that the defect density remains extensive up to a time $T_{\text{plateau}}\sim L$; then, a rapid decay takes place, with $n\sim T^{-2}$. It follows that the onset of adiabaticity, defined by $n(T_{\mathrm{ons}})\approx 1/L$, takes a time $T_{\mathrm{ons}}\sim L^{3/2}$. Thus, the QAB approach does not provide efficient schedules $g_{\mathrm{QAB}}(t)$, as it is still governed by the local energy gap. Even though the zero mode can be driven adiabatically in a timescale $T\sim L$, this comes at the price of creating an extensive number of excitations at the avoided level crossing of higher momentum modes. Similarly to the LAD approach, the universal KZ regime does not emerge during the whole process.
@@ IV Optimal control beyond the adiabatic condition
In Sec. III, we considered optimal control strategies that take as starting point the adiabatic condition, i.e., that require the drive to be slow with respect to the inverse gap with (Eq. (15)) or without (Eq. (24)) the transition matrix elements. In this Section, we introduce optimal control methods that go beyond the adiabatic condition, taking into account also the higher excited states. Specifically, in Sec. IV.1, we extend the QAB to tame also transitions in excited states, while in Sec. IV.2, we introduce a new action to be minimized, based on the avoided level crossing structure in the LZ formula. Finally, we compare our results with other well-known optimization schemes: the ONLP through a quantum critical point in Sec. IV.3 and the fully numerical CRAB optimization in Sec. IV.4.
@@ IV.1 Extended quantum adiabatic brachistrochrone
where |$\ket{\epsilon_{\pm,k}(t)}$ denotes the ground (--) and excited ($+$) single-particle states in the $k$-th momentum mode, and $2\epsilon_{k}(t)$ equals the gap in the same mode. We call this action “extended quantum adiabatic brachistochrone” (eQAB). Notice that in the numerator of Eq. (34) there is the matrix element between the one-particle ground and excited states, rather than the norm $\|\partial_{t}\hat{H}\|$, as described in Sec. III.3. This choice allows for more accurate handling of the transition rates in the independent TLSs.
@@
The reason why the eQAB approach does not provide a significant improvement over the QAB can be understood as follows. Equation (34) is an unweighted average over all momentum modes of an action of the QAB form. There are two consequences to this fact. First, the zero-momentum mode still dominates, as exemplified by the Euler-Lagrange equation becoming singular in the thermodynamic limit. Second, the main contribution to the excitations arises from the vanishingly small regions around the avoided crossings. However, the action in Eq. (34) does not weight properly these short time intervals where the gaps are closing, masking their true impact. In addition, the non-adiabatic rates within the eQAB action do not account accurately for excitation probabilities around the avoided crossings. In other words, the LZ transitions are insensitive to the local adiabatic conditions when the latter is shorter than the global adiabatic time-scale, $T<T_{\mathrm{ad}}$. Increasing the time intervals where the local adiabatic conditions are taken into account misses the source of non-adiabaticity. These two facts lead to a considerably weaker effective condition to suppress excitations.
@@ IV.2 Non-Adiabatic Quantum Optimization (NAQO) of Landau-Zener transitions
Above, it was argued that tailoring control schedules by the local adiabatic condition, Eq. (14), leads to suboptimal results, as higher-excited modes are simply not considered. Remarkably, no relevant improvement in the timescale of the onset of adiabaticity was found even by the simultaneous optimization of all the momentum modes with the eQAB, see Sec. IV.1. These results indicate that even in simple cases, many-body effects lead to non-trivial behavior, deteriorating the efficiency of the conventional optimal control strategies.
In this section, we show that it is possible to design better schedules by focusing attention on the main source of non-adiabatic effects, i.e., the small neighborhood of the avoided level crossings. As anticipated at the end of the last section, the schedule $g(t)$ needs to go slow only when there is a high probability of exciting a momentum mode, i.e., near an avoided crossing. This is exactly where the LZ formula applies, both in the KZ scaling regime and beyond the onset of adiabaticity [12, 78, 82, 83, 13]. Designing efficient control schedules requires the simultaneous suppression of LZ transitions in both low and high-momentum states. To capture this in an analytical way, the following action is introduced:
@@
where $g_{c}(k):=g(t_{c}(k))$ is the position of the avoided crossing as a function of $k$; see Eq. (10). We denote Eq. (41) as the NAQO action. This action can also be understood as refining the uniform averaging over time in Eq. (24) to only those time instances where the $k$-th avoided crossing is formed, fixed by the relation $g(t^{*})=\cos k$, Eq. (10). Additionally, the integrand is modified by using the LZ formula instead of the local-adiabatic transition rate. Notice that $S_{\mathrm{NAQO}}$ is defined only for values of $g<1$, i.e., where there actually are avoided crossings. How to extend the action for $g>1$ is explained below.
@@ IV.4 Fully numerical optimization
Finally, let us illustrate the shape of the considered optimal control protocols in Fig. 7. Interestingly, the typical shape of the extended QAB schedules is of the opposite nature to those of the LAD and QAB approaches. These latter are extremely flat around $g=1$, with a slope scaling as $\sim(TL)^{-1}$, and present a sharp jump around $t=\pm T/2$. The former is more similar to the linear schedule, without dramatic changes in the local velocity $\dot{g}(t)$. The ONLP and NAQO protocols exhibit features are of an intermediate structure, as they tend to change slowly around the critical point, however do not speed up significantly at high momenta either. However, the NAQO schedule exhibits a more optimal shape balancing the simultaneous effects of both low and high momentum states. In contrast, the CRAB approach exhibits additional features and a surprising non-monotonic behavior close to the QPT.
Figure 6: Comparison of the analytical NAQO and ONLP with the average and best CRAB numerical optimization. The system size is fixed with $L=1000$ since the CRAB algorithm becomes numerically demanding for larger systems. While, on average, the CRAB search does not find good schedules, the best instance over $1000$ trials falls near the ONLP results. The NAQO results show an anomalously fast decay here due to finite-size effects, while they approach $n\sim T^{-1}$ at larger sizes, see Fig. 4.
@@ V.2 Optimal control in the universal KZ scaling regime
In this regime, the behavior of the low momentum modes slightly changes, and a more pronounced role goes to their suppression, as entailed by $\dot{g}^{3}/(2-g)^{(2\beta+\alpha-3)/(\alpha-1)}$. The second term, however, only acquires an $\alpha$ and $\beta$ dependent prefactor, accounting in the same way for the higher momentum excitations. This feature is attributed to the additional $\alpha$-dependent exponent in the corresponding differential equation. Despite these additional features, the resulting NAQO schedule provides the fastest decay of the defects for all $\alpha<3$.
@@
where again $W$ is the Lambert function, cf. Eq. (46). At small $t$ values, the denominator can be neglected, while for intermediate values, the shape stays close to the linear one as well. Note that this limiting differential equation is of completely opposite nature to that in the eQAB, Eq. (40), which only involves the $\dot{g}^{2}/(2-g)$ term.
Finally, in the case of Eq. (72), the $\alpha$ independence is only observed for the linear and ONLP schedules, while increasing $\alpha$ shifts the LZ-optimized schedules by a constant factor, as observed in 12b,c.
@@ V.3 Optimal control in the dynamical power-law-scaling regime
where notably the low-momentum term on the r.h.s., $\dot{g}^{3}/(1-g)^{\frac{\alpha+1}{\alpha-1}}$, only depends now on $\alpha$. In this regime, both ONLP and NAQO provide a faster decay of $n(T)$ as $\alpha$ increases, while the linear schedule results remain unchanged according to Eq. (75), as predicted in Ref. [82].
@@ VI Demonstration on a non-integrable disordered many-body system
The model in Eq. (78) does not admit a separation into independent momentum modes, but instead, the Hilbert space is composed of only two $2^{L-1}$-dimensional sectors, according to the parity $\prod_{i}\hat{\sigma}_{i}^{x}$. We concentrate on the even parity sector, which is the one dynamically accessible from the paramagnetic ground state.
As in the integrable cases, excitations are generated mainly via the LZ transitions at multiple avoided crossings. Although these transitions are no longer independent of each other, and their analytical treatment is not possible, some robust features are still present. In particular, avoided crossings of higher excited states appear with gradually increasing gaps and at smaller values of $g(t)$, motivating the framework of the NAQO. As a stepping stone towards a more precise numerical treatment of these avoided crossings, we employ an ansatz approach in terms of the general LRKM control equation. Even though the critical exponents are not accessible analytically, within this ansatz approach, we identify the $\beta$ parameter by the $L$ dependence of the ground state gap. For the numerically achievable system sizes $L\lesssim 20$, the ground-state energy gap approximately decreases as $\sim L^{-0.65}$ (see inset of Fig. 15). The general LRKM equation in long-range hopping and short-range pairing regime $\beta<2,\,\alpha>3$ (see Sec. V) reads
@@
In the first term on the r.h.s., the $L$ dependence of the gap at the first avoided level crossing is captured by $\beta$. The second term accounts for the suppression of LZ transitions to higher excited states. For the numerically reachable system sizes, $L\lesssim 20$, the effective system size dependence of the ground state gap sets the exponent based on the dispersion relation of the LRMs, $\beta=1.65$, while the second term on the r.h.s. provides a good approximation for the higher excited states as it depends on $\beta$ only through a multiplicative factor. Additionally, the limit $\alpha\rightarrow\infty$ was taken to illustrate the use of this ansatz in the simplest possible form. The initial condition of the corresponding optimal differential equation is identified with the position of the gap to the ground state, also found numerically. Thus, the corresponding optimal control equation reads
@@
The results are shown in Fig. 15 for $L=14$ nd $L=18$. The present method provides a tangible speed-up compared to earlier path optimization methods [108, 109, 110], which rely exclusively on numerical optimization and typically require also the use of intermediate control Hamiltonians. In these studies, both the computational resources and the achievable efficiency are much larger than the NAQO method either semi-analytical in integrable models or assisted by additional numerical techniques in non-integrable ones.
@@ VII Conclusions and perspectives
A challenging and promising prospect is to apply NAQO and the LZ formula to non-integrable models and in more general quantum annealing contexts. While we have demonstrated for the specific example of Sec. VI that NAQO does provide some speed-up at small system sizes, it is an open question whether NAQO is able to speed up the annealing also when, at larger sizes, the scaling of the gap becomes exponential in the system size [12]. The same exponential scaling is at the base of the difficulty of applying quantum annealing to NP-hard problems [105, 106]. In this context, optimal control is still largely uncharted, and the few existing studies point to the presence of interesting phases and controllability transitions [111, 112, 113, 114]. Another interesting direction would be to use the analytically-optimized NAQO schedules as a starting point for further numerical optimization in the spirit of Refs. [58, 115].